### v1

@@ Abstract
We study decision making under risk when perception may be menu-dependent. Behavior is modeled as the outcome of a small library of psychologically interpretable, menu-specific rules that transform each objective menu into a perceived one. At each menu, the applied rule must make the realized choice a strict improvement under a dominance benchmark on perceived lotteries. We introduce the Maximum Rule Concentration Index (MRCI), the maximal Herfindahl-Hirschman concentration of rule shares over all locally admissible assignments, and diagnostics that distinguish rules that unify behavior across many menus from rules that mainly act as substitutes. We provide a MIQP formulation, a scalable heuristic, and a finite-sample permutation test of excess concentration relative to a menu-independent random-choice benchmark. Applied to the CPC18 dataset ($N=686$ subjects, each making 500500-$700$ repeated binary lottery choices), the mean MRCI is $0.545$, and $64.1\%$ of subjects reject random choice at the $1\%$ level. Concentration gains are primarily driven by modal-payoff focusing, salience-thinking, and regret-based comparisons.
@@ 1 Introduction
Choices under risk are famously patterned. In laboratories and in the field, people violate monotonicity, reverse rankings when lotteries are mixed with a common consequence, and display menu-dependent behavior that resists any single, stable ordering over objective lotteries. Over the past decades, the literature has responded with an impressive array of models - probability weighting, reference dependence, salience, regret, disappointment, limited attention - each offering a psychologically plausible channel through which the same objective menu may be perceived and evaluated differently. [^1] See, among many others, kahneman_tversky1979; tversky_kahneman1992; prelec1998 on probability weighting; KoszegiRabin2006 on reference dependence; bordalo2012_risk on salience in risky choice; LoomesSugden1986; bell85; Gul1991 on regret and disappointment; and sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014 on limited attention and consideration. This richness has greatly expanded our understanding of risky choice, but it also raises a complementary empirical question. When perception may vary with the menu, many distinct behavioral mechanisms can potentially explain the same set of risky decisions. This motivates a different emphasis: in a large dataset, which behavioral mechanisms are actually carrying the explanatory burden, how many distinct mechanisms are needed to organize behavior, and how stable those conclusions are as the set of candidate explanations is enlarged.
In this paper, we study binary risky choice through the lens of menu-specific decision rules. We posit a finite library of psychologically interpretable transformations - worst-case focusing (max-min), coarse probability distortion, salience-based focusing on the most contrasting state, regret- and disappointment-based comparisons, and one-sided attention that map each objective binary menu into a perceived one. The analyst does not observe these perceptions; she observes only the realized choices. Our identifying restriction is therefore local: at each menu, the rule assigned to that menu must render the chosen perceived lottery a strict improvement under a First Order Stochastic Dominance. This menu-level discipline has a sharp consequence: any locally admissible rule profile induces a perceived dataset that is automatically cycle-free in revealed-preference terms. Thus, the empirical content of the framework is not obtained by imposing global transitivity on observed choices. It comes from asking how parsimoniously one can assign rules across menus while preserving local dominance-based justification.
@@
We also develop inference. High concentration might be mechanically attainable by chance, even in environments where the admissibility constraints are nontrivial. To address this, we propose a finite-sample permutation test in the spirit of cherchye2025. The test is grounded in a Random Rule Model interpretation that is specific to binary menus. Because each choice problem has two labeled alternatives (left/right) and the library includes the two one-sided attention defaults (“attend to the left” and “attend to the right”), any realized sequence of left/right choices can be replicated by random switching between these two rules. [^2] This is similar to treatments of the power of nonparametric tests of preference maximization.See bronars and beatty2011. Under the null that choices are exchangeable across menus (menu-independent randomness), the conditional distribution of the MRCI given the realized number of left versus right choices is obtained by permuting the left/right indicators across fixed menus. We compare the observed MRCI to the conditional randomization distribution. Rejection means that the observed maximal concentration is too large to be plausibly generated by menu-independent random switching, providing statistical content to the notion that “few-rule behavior” is genuinely present in the data.
@@
Our inference procedure shows that this concentration is not merely mechanical. Implementing the permutation test, we find that $64.1\%$ of subjects reject the random-choice benchmark at the $1\%$ level. In other words, for a large share of subjects, the observed degree of concentration is too high to be plausibly generated by a menu-independent random process once we condition on the realized frequency of left/right choices.
@@
The rule-importance diagnostics then identify which mechanisms account for this concentration. We find that MAP (modal-payoff focusing) is the main unifier on average, accounting for about $17\%$ of attainable concentration, with SAL (salience-based focusing) next at roughly $7\%$; regret-based comparisons (REG) and the extremum heuristics (MMx/MMn) contribute smaller but meaningful shares (on the order of $4$-$5\%$ each). Stability is more evenly distributed across the active rules: while MAP is most often indispensable (with a stability score $40\%$), REG and SAL also remain necessary in a nontrivial fraction of cases (roughly $25\%$ and $20\%$), indicating substantial substitution across mechanisms even when MAP drives the average gain. By contrast, PW and Id have negligible importance. This might be because these rules often fail to induce strict FSD improvements menu-by-menu. DIS appears genuinely non-operative in CPC18, in the sense that it almost never affects maximal concentration and is rarely needed to sustain it.
@@
Our framework speaks to three strands of the literature. First, it complements the large decision-theoretic literature that explains patterned risky choice by enriching preferences or perception, including probability weighting, reference dependence, salience, regret, disappointment, and limited attention (kahneman_tversky1979; tversky_kahneman1992; prelec1998; KoszegiRabin2006; bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; masatlioglu2012; manzini_mariotti2014; caplin_dean2015). We complement this literature, as we provide a nonparametric way to summarize how much structure is needed to render behavior locally defensible, and which broad mechanisms are most consistent with the data once they compete on a common footing. This perspective differs from approaches that prioritize flexible prediction with large, high-dimensional models or machine-learning toolkits, including recent work that achieves high predictive accuracy in risky-choice prediction settings, including erev2011, erev_ert_plonsky_2017, plonsky2016, or peterson2021.
@@
Second, the paper connects to the growing literature that models choice through procedures and simple rules rather than through a single stable preference ordering. This tradition includes bounded rationality and heuristic choice (simon1955; tversky_kahneman1974; Kahneman1982; rubinstein88), “rule rationality” arguments emphasizing the efficiency of simple rules in natural environments and their fragility in contrived ones (aumann2008; aumann2019), simplicity in choices under risk (enke2023; fudenberg2022; puri2025), automata-based complexity constraints (salant2011menu), and procedural models of consideration, shortlisting, and categorization (salant2006; salant_rubinstein2008; manzini2007; manzini2012; cherepano2013), as well as recent experimental work that studies procedure choice directly (halevy2024). We are deliberately agnostic about why a given decision maker uses a particular heuristic or when a rule is optimal; our focus is on identification in standard risky-choice datasets, where rules are latent and must be inferred from choices.
@@
Third, the paper contributes to the revealed preference literature. Classic revealed-preference theory, from samuelson1938note onward, makes cyclical consistency central by treating global rationalizability as the core testable implication of utility maximization (afriat1967; varian1982; houthakker1950). [^3] Using an experimental design that extends the canonical budget-line paradigm of syngjoo2007 to a three-dimensional risky-choice environment, dembo2021 find that departures from expected-utility theory go beyond violations of independence, and that many prominent non-EUT alternatives - which nonetheless preserve first-order stochastic dominance - do not fully account for these patterns. polisson2020 develops a nonparametric revealed-preference approach for testing a broad class of choice under risk models using data from budget-line experiments. In our framework, discipline comes instead from a local dominance-based justification applied after rule-induced transformations. Under first-order stochastic dominance, this local discipline implies a strong form of cyclical consistency in the induced perceived dataset. Methodologically, we also complement the revealed-preference computation and inference literature that relies on mixed-integer programming to compute indices and test statistics (HEUFER201587; Demuynck2023) by providing a scalable heuristic for our concentration object, together with a permutation test similar to cherchye2025.
@@
The rest of the paper proceeds as follows. Section 2 introduces the choice environment, the rule library, and formalizes admissible rule profiles. Section 3 develop several examples, and Section 4 defines the MRCI and the rule-importance diagnostics. Section 5 develops a scalable computational procedure. Section 6 provides the permutation test. Section 7 reports the empirical results, and Section 8 concludes.
@@ 2.1 Bounded Rationality
The rule library is deliberately simple and nonparametric. It collects ten parameter-free rules that span four broad mechanisms emphasized in the bounded-rationality and behavioral-decision literatures: outcome simplification, probability distortion, context-dependent comparisons, and limited attention or consideration. [^4] On simple heuristics and “fast and frugal” rules see, e.g., tversky_kahneman1974; gigerenzer_gaissmaier2011; brandstatter_et_al2006. On probability weighting and cumulative prospect theory, see kahneman_tversky1979; tversky_kahneman1992; prelec1998. On salience, regret and disappointment, see bordalo2012_risk, LoomesSugden1986; bell85; Gul1991, and KoszegiRabin2006. On limited attention and consideration, see, among others, sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015. We now define each rule formally.
@@
The library is deliberately simple and nonparametric: all ten rules are parameter-free and are designed as coarse, reduced-form implementations of four broad mechanisms central in the bounded-rationality and behavioral-decision literatures. MMn, MMx and MAP implement outcome-simplification heuristics (rubinstein88; gigerenzer_gaissmaier2011; brandstatter_et_al2006) by projecting each lottery onto a single payoff attribute (minimum, maximum, or modal payoff). PW distorts probabilities in a coarse fashion that echoes probability weighting in cumulative prospect theory (kahneman_tversky1979; tversky_kahneman1992; prelec1998). SAL, REG and DIS capture different forms of context-dependent comparisons across and within lotteries, providing nonparametric counterparts to salience, regret and disappointment models (bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; KoszegiRabin2006). Finally, A1 and A2 encode extreme one-sided attention and thus serve as simple menu-level representations of limited attention and consideration-set formation (sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015). All ten rules fit the general menu-transformer framework $f(L^{1},L^{2})=(\pi_{f}(L^{1};L^{2}),\pi_{f}(L^{2};L^{1}))$. The rule library introduced in this section is tailored to binary lottery choice environments. There might be additional relevant rules for more complex choice environments, including sampling-based perception, support compression, or rank-dependent distortions beyond a single probability. [^6] See, for example, hu2023. The next subsection specifies when a profile of such rules is considered an admissible explanation of a dataset, and how we quantify the parsimony of these explanations.

### v2

@@ Abstract
We study choice among lotteries in which the decision maker chooses from a small library of decision rules. At each menu, the applied rule must make the realized choice a strict improvement under a dominance benchmark on perceived lotteries. We characterize the maximal Herfindahl-Hirschman concentration of rule shares over all locally admissible assignments, and diagnostics that distinguish rules that unify behavior across many menus from rules that mainly act as substitutes. We provide a MIQP formulation, a scalable heuristic, and a finite-sample permutation test of excess concentration relative to a menu-independent random-choice benchmark. Applied to the CPC18 dataset ($N=686$ subjects, each making 500500-$700$ repeated binary lottery choices), the mean rule concentration is $0.545$, and $64.1\%$ of subjects reject random choice at the $1\%$ level. Concentration gains are primarily driven by modal-payoff focusing, salience-thinking, and regret-based comparisons.
@@ 1 Introduction
Choices under risk are famously patterned. In laboratories and in the field, people violate monotonicity, reverse rankings when lotteries are mixed with a common consequence, and display menu-dependent behavior that resists any single, stable ordering over objective lotteries. Over the past decades, the literature has responded with an impressive array of models - probability weighting, reference dependence, salience, regret, disappointment, attenuation, cognitive uncertainty, limited attention - each offering a psychologically plausible channel through which the same objective menu may be perceived and evaluated differently. [^1] See, among many others, kahneman_tversky1979; tversky_kahneman1992; prelec1998 on probability weighting; KoszegiRabin2006 on reference dependence; bordalo2012_risk on salience in risky choice; LoomesSugden1986; bell85; Gul1991 on regret and disappointment; enke2024_attenuation on behavioral attenuation, enke2023_uncertainty on cognitive uncertainty, and sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014 on limited attention and consideration. This richness has greatly expanded our understanding of risky choice, but it also raises a complementary empirical question. When perception may vary with the menu, many distinct behavioral mechanisms can potentially explain the same set of risky decisions. This motivates a different emphasis: in a large dataset, which behavioral mechanisms are actually carrying the explanatory burden, how many distinct mechanisms are needed to organize behavior, and how stable those conclusions are as the set of candidate explanations is enlarged.
In this paper, we study binary risky choice through the lens of menu-specific decision rules. We posit a finite library of psychologically interpretable transformations - worst-case focusing (max-min), coarse probability distortion, salience-based focusing on the most contrasting state, regret- and disappointment-based comparisons, and one-sided attention that map each objective binary menu into a perceived one. The analyst does not observe these perceptions; she observes only the realized choices. Our identifying restriction is therefore local: at each menu, the rule assigned to that menu must render the chosen perceived lottery a strict improvement under First Order Stochastic Dominance. Thus, the empirical content of the framework is not obtained by imposing global transitivity on observed choices. It comes from asking how parsimoniously one can assign rules across menus while preserving local dominance-based justification.
@@
We also develop inference. High concentration might be mechanically attainable by chance, even in environments where the admissibility constraints are nontrivial. To address this, we propose a finite-sample permutation test in the spirit of cherchye2025. The test is grounded in a Random Rule Model interpretation that is specific to binary menus. Because each choice problem has two labeled alternatives (left/right) and the library includes the two one-sided attention defaults (“attend to the left” and “attend to the right”), any realized sequence of left/right choices can be replicated by random switching between these two rules. [^2] This is similar to treatments of the power of nonparametric tests of preference maximization. See bronars and beatty2011. Under the null that choices are exchangeable across menus (menu-independent randomness), the conditional distribution of the MRCI given the realized number of left versus right choices is obtained by permuting the left/right indicators across fixed menus. We compare the observed MRCI to the conditional randomization distribution. Rejection means that the observed maximal concentration is too large to be plausibly generated by menu-independent random switching, providing statistical content to the notion that “few-rule behavior” might be present in the data. We show that the statistical test has asymptotic power of $1$. Intuitively, when there is genuine excess latent concentration, the realized MRCI eventually lies far in the upper tail of the conditional randomization distribution.
@@
Our inference procedure shows that this concentration is not merely mechanical. Implementing the statistical test with $500$ permutations, we find that $64.1\%$ of subjects reject the random-choice benchmark at the $1\%$ level. In other words, for a large share of subjects, the observed degree of concentration is too high to be plausibly generated by a menu-independent random process once we condition on the realized frequency of left/right choices.
@@
The rule-importance diagnostics then identify which mechanisms account for this concentration. We find that MAP (modal-payoff focusing) is the main unifier on average, accounting for about $17\%$ of attainable concentration, with SAL (salience-based focusing) next at roughly $7\%$; regret-based comparisons (REG) and the extremum heuristics that compare either the maximum or the minimum payoffs of the two lotteries (MMx/MMn) contribute smaller but meaningful shares (on the order of $4$-$5\%$ each). Stability is more evenly distributed across the active rules: while MAP is most often indispensable (with a stability score $40\%$), REG and SAL also remain necessary in a nontrivial fraction of cases (roughly $25\%$ and $20\%$), indicating substantial substitution across mechanisms even when MAP drives the average gain.
By contrast, PW (Probability Weighting) and Id (Identity) have negligible importance by our two rule-importance diagnostics. This does not mean that these transformations are rarely locally admissible: in CPC18, both rules justify the realized choice under the strict FSD discipline for a nontrivial share of menus. Rather, their low concentration gains and low stability indicate that they are seldom pivotal for attaining maximal concentration once other mechanisms are available: menus that can be assigned to Id or PW can typically be reassigned to the dominant “unifiers” (notably MAP and SAL) with little loss in MRCI. DIS (disappointment) is similar: while it can be locally admissible at some menus, it almost never becomes necessary to sustain a high-concentration assignment, suggesting that disappointment-style comparisons may not be a leading organizer of CPC18 behavior.
@@
Our framework speaks to three strands of the literature. First, it complements the large behavioral literature that explains patterned risky choice by enriching preferences or perception, including probability weighting, reference dependence, salience, regret, disappointment, cognitive uncertainty, and limited attention (kahneman_tversky1979; tversky_kahneman1992; prelec1998; KoszegiRabin2006; bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; masatlioglu2012; manzini_mariotti2014; caplin_dean2015; enke2023_uncertainty). Our approach is deliberately nonstructural relative to these models. Rather than positing a utility representation - either in the objective domain or in an agent’s perceived domain - we impose a purely local discipline: at each menu, the rule assigned to that menu must transform the objective pair into a perceived pair in which the realized choice is a strict improvement under a dominance benchmark. We then ask how parsimoniously such locally defensible explanations can be organized across menus, and which broad mechanisms most often survive when they compete on a common footing. This perspective also differs from approaches that prioritize flexible prediction with large, high-dimensional models or machine-learning toolkits, including recent work that achieves high predictive accuracy in risky-choice prediction settings (erev2011; erev_ert_plonsky_2017; plonsky2016; peterson2021).
@@
Second, the paper connects to the growing literature that models choice through procedures and decision rules. This tradition includes bounded rationality and heuristic choice (simon1955; tversky_kahneman1974; Kahneman1982; rubinstein88), “rule rationality” arguments emphasizing the efficiency of simple rules in natural environments and their fragility in contrived ones (aumann2008; aumann2019), simplicity in choices under risk (enke2023; fudenberg2022; puri2025; oprea2024; hu2023), automata-based complexity constraints (salant2011menu), and procedural models of consideration, shortlisting, and categorization (salant2006; salant_rubinstein2008; manzini2007; manzini2012; cherepano2013), as well as recent experimental work that studies procedure choice directly (arrieta2024; halevy2024). We are deliberately agnostic about why a given decision maker uses a particular heuristic or when a rule is optimal; our focus is on identification in standard risky-choice datasets, where rules are latent and must be inferred from choices.
@@
Third, the paper contributes to the revealed preference literature. Classic revealed-preference theory, from samuelson1938note onward, makes cyclical consistency central by treating global rationalizability as the core testable condition for utility maximization (afriat1967; varian1982; houthakker1950). [^3] Using an experimental design that extends the canonical budget-line paradigm of syngjoo2007 to a three-dimensional risky-choice environment, dembo2021 find that departures from expected-utility theory go beyond violations of independence, and that many prominent non-EUT alternatives - which nonetheless preserve first-order stochastic dominance - do not fully account for these patterns. polisson2020 develops a nonparametric revealed-preference approach for testing a broad class of choice under risk models. In our framework, discipline comes instead from a local dominance-based justification applied after rule-induced transformations. Under first-order stochastic dominance, we show that this local discipline implies a strong form of cyclical consistency in the induced perceived dataset. Methodologically, we also complement the revealed-preference computation and inference literature that relies on mixed-integer programming to compute indices and test statistics (HEUFER201587; Demuynck2023) by providing a scalable heuristic for our concentration object, together with a permutation test similar to cherchye2025.
@@
The rest of the paper proceeds as follows. Section 2 introduces the choice environment, the rule library, and formalizes admissible rule profiles. Section 3.3 develop several examples, and Section 3 defines the MRCI and the rule-importance diagnostics. Section C develops a scalable computational procedure. Section 4 provides the permutation test. Section 5 reports the empirical results, and Section 6 concludes.
@@ 2.1 Bounded Rationality
The rule library is deliberately simple and nonparametric. It collects ten parameter-free rules that span four broad mechanisms emphasized in the bounded-rationality and behavioral literatures: outcome simplification, probability distortion, context-dependent comparisons, and limited attention or consideration. [^4] On simple heuristics and “fast and frugal” rules see, e.g., tversky_kahneman1974; gigerenzer_gaissmaier2011; brandstatter_et_al2006. On probability weighting and cumulative prospect theory, see kahneman_tversky1979; tversky_kahneman1992; prelec1998. On salience, regret and disappointment, see bordalo2012_risk, LoomesSugden1986; bell85; Gul1991, and KoszegiRabin2006. On limited attention and consideration, see, among others, sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015. We now define each rule formally.
@@
The library is deliberately simple and nonparametric: all ten rules are parameter-free and are designed as coarse, reduced-form implementations of four broad mechanisms central in the bounded-rationality and behavioral-decision literatures. MMn, MMx and MAP implement outcome-simplification heuristics (rubinstein88; gigerenzer_gaissmaier2011; brandstatter_et_al2006) by projecting each lottery onto a single payoff attribute (minimum, maximum, or modal payoff). PW distorts probabilities in a coarse fashion that echoes probability weighting in cumulative prospect theory (kahneman_tversky1979; tversky_kahneman1992; prelec1998). SAL, REG and DIS capture different forms of context-dependent comparisons across and within lotteries, providing nonparametric counterparts to salience, regret and disappointment models (bordalo2012_risk; LoomesSugden1986; bell85; Gul1991; KoszegiRabin2006). Finally, A1 and A2 encode extreme one-sided attention and thus serve as simple menu-level representations of limited attention and consideration-set formation (sims2003; gabaix2014; masatlioglu2012; manzini_mariotti2014; caplin_dean2015). All ten rules fit the general menu-transformer framework $f(L^{1},L^{2})=(\pi_{f}(L^{1};L^{2}),\pi_{f}(L^{2};L^{1}))$. The rule library introduced in this section is tailored to binary lottery choice environments. There might be additional relevant rules for more complex choice environments, including sampling-based perception, support compression, or rank-dependent distortions beyond a single probability. [^6] See, for example, hu2023.
Library choice, coarseness, and interpretation.
Our objects are defined conditional on a candidate library $\mathcal{F}$: the admissible profile set, the induced rule shares, and the resulting concentration measures all depend on which menu transformations are allowed. This dependence is not a drawback but a design choice. The goal is not to estimate a single structural model of risky choice; it is to ask how parsimoniously a dataset can be organized within a transparent set of broad mechanisms.
The library is intentionally coarse and parameter-free. A natural concern is that some rules might be mechanically irrelevant because they rarely pass the strict admissibility screen. To address this, we report a rule-by-rule coverage statistic in the application: for each subject and each rule $f\in\mathcal{F}$, coverage is the fraction of that subject’s observations for which $f$ is strictly admissible (i.e., yields a FSD improvement relative to the realized choice). We then average this quantity across subjects. As shown by the coverage figure, all rules in $\mathcal{F}$ exhibit substantial coverage in the data, so low estimated importance for a rule reflects substitutability or lack of pivotal explanatory power rather than failure to ever qualify as admissible.
Additionally, as demonstrated later on, enlarging the rule library can only weakly increase maximal concentration. In this sense, our baseline library yields a conservative benchmark for “how concentrated” behavior can look under a standard, interpretable set of mechanisms.
Alternative libraries and nesting.
The framework is modular, as any finite collection of menu transformations can be used in place of our baseline $\mathcal{F}$, and all definitions and computations carry over verbatim. One natural robustness exercise is to consider nested libraries $\mathcal{F}_{1}\subseteq\mathcal{F}_{2}\subseteq\cdots$ that progressively enrich the set of candidate mechanisms. In principle, this would let one trace a “concentration–richness” curve and ask how quickly maximal concentration increases as additional forms of context dependence are allowed. In the application section, we build two metrics in this spirit.
In practice, constructing a nesting that simultaneously and fairly “contains” the main behavioral theories is challenging. Many theories correspond to parameterized families of perception or evaluation maps, and embedding each family into a common, comparable sequence may require choosing functional forms, discretization grids, and relative weights across families. Different choices can change the effective richness of each mechanism and therefore tilt the comparison in ways that are hard to justify a priori. Moreover, mechanisms such as salience and regret are qualitatively distinct rather than ordered by inclusion. For these reasons, we adopt a transparent, theory-driven baseline library of coarse, interpretable rules. Generating rules from common primitives is an important direction for future work.